Last year Sue was three times as old as Pablo. Next year she will be twice as old as Pablo is. How old is each now?
I don't understand this at all.): Could someone please help me?

Answer by ankor@dixie-net.com(22740)

Last year Sue was three times as old as Pablo. Next year she will be twice as old as Pablo is. How old is each now?
:
This is not hard
:
Let s = Sue's age now, Let p = Pab's age now
:
Last year their ages would be:
(s-1) and (p-1), right?
:
It said,"Last year Sue was three times as old as Pablo." so we can say:
(s-1) = 3(p-1)
s - 1 = 3p - 3; just do some basic algebra on it
s = 3p - 3 + 1
s = (3p - 2)
:
Next year she will be twice as old as Pablo is. How old is each now?
:
Next year their ages would be:
(s+1) and (p+1), right?
:
It says,"Next year she will be twice as old as Pablo is." so we can say:
(s+1) = 2(p+1)
s + 1 = 2p + 2
s = 2p + 2 - 1
s = 2p + 1
:
How old is each now?:
Using the 1st equation we found s = (3p-2), substitute that for s in the 2nd eq
3p - 2 = 2p + 1
3p = 2p + 1 + 2
3p - 2p = 3
p = 3 yrs is Pab's present age
:
Find Sue's age using s = 2p + 1
s = 2(3) + 1
s = 6 + 1
s = 7 yrs is Sue's present age
:
:
Check out our solutions by using the two sentences:
Last year Sue was three times as old as Pablo.
(7-1) = 3(3-1) true
and
Next year she will be twice as old as Pablo is.
(7 + 1) = 2(3+1) also true
:
:
A lot of steps, but each step is logical and simple, did you understand this?